Pens are sold in bundles of 25. For every 5 bundles of pens you will get an extra of 3 pens. If you want to buy 364 pens for your friends, how many bundles of pens you must buy?
Answer by richwmiller(17219)

Trick question!
Do you mean buy or obtain?
Do you want to include the free pens?
As the question is written, the answer is no.
As it turns out, the question is moot.
364/25=14.56 bundles
14*25=350
You receive 6 pens for the 10 bundles.
350+6=356
364-356=8
You are still short 8 pens.
So to reach your goal of 364 pens, you must buy another bundle for a total of 15 bundles.
15*25=375
You will receive 15/5=3
3*3=9 free pens
You will receive 375 +9=384 pens
Conclusion:
You must buy a total of 15 bundles.